5. You have some savings you’d like to invest. You can invest it in a bank account paying fixed 5% compounded monthly. Your plan is to leave the principal but take out the interest at the end of every year to help cover expenses. If you wanted to be sure that your savings invested at the bank allowed you to take out $5,000 per year for as long as you wanted to, how much savings would you have to have?
Answer #1

EAR = (1+APR)^n - 1

= (1+0.05/12)^12 - 1

= 0.05116189788

present value of perpetuity = cash flow per year/interest rate

= 5000/0.05116189788

= 97728.98 ...........ans